Find the producer's surplus at Q = 9 for the following supply functions:



(a) P = 12 + 2Q



(b) P = 20√Q + 15

1
Expert's answer
2021-11-25T20:53:01-0500

a) ="P=12+2Q"

At Q=9,P=12+18=30


Therefore producer's surplus ="P. Q-\int PdQ"


"=(30)(9)-\int(12+2Q)dQ"


"=270-[12Q+Q^2] ^9_0"


"=270-[108+81]"


"=81"


b) "P=20\sqrt{Q}+15"


At Q=9,"P=20\sqrt{9}+15 =75"


Producer's Surplus ="PQ-\int PdQ"


"=(75)(9)-\int 20\sqrt{Q}+15 dQ"


"=675-[20\frac {Q^{\frac {3} {2} }}{\frac {3} {2}} +15Q]^9_0"


"=675-495"


"=180"